E ⊗ e VIBRONIC COUPLING IN QUARTET STATES OF Rb_{3} AND K_{3} ON HELIUM DROPLETS

GERALD AUBÖCK, JOHANN NAGL, \underline{CARLO CALLEGARI}, and WOLFGANG E. ERNST, Institute of Experimental Physics, TU Graz, Petersgasse 16, 8010 Graz, Austria/EU.

We present laser-induced-fluorescence (LIF) and magnetic-circular-dichroism (MCD) spectra of the (2)^{4}E′ ← (1)^{4}A′_{2} band of Rb_{3} and K_{3} on He nanodroplets. For Rb_{3} we find a progression of four bands, split by spin-orbit (SO) and weakly perturbed by linear vibronic Jahn-Teller (JT) coupling. For K_{3} the transition had been previously observed, and interpreted in terms of a linear and quadratic Jahn-Teller coupling in the (2)^{4}E′ state.^{a} An unambiguous assignment of spectra on He droplets is however difficult due to matrix broadening; MCD spectra are a valuable method to offset this unwanted effect. For K_{3} they clearly reveal the importance of SO coupling and suggest a different assignment with weak linear, and no significant quadratic, JT coupling. Further, interpretation of our data in terms of free-molecule SO indicate that the trimers lie flat on the droplet surface. The observed strong C-type MCD spectra arise from different populations of the ground-state Zeeman sublevels; a quantitative analysis allows a determination of the spin temperature, hence of the temperature at the droplet surface. Our simulations are consistent with a value of 0.4 K, which agrees with our previous findings from spectra of triplet-state alkali-metal dimers.^{b} This is the first observation of SO coupling in quartet states of trimers.

aJ. H. Reho, J. Higgins, M. Nooijen, K. K. Lehmann, and G. Scoles, J. Chem. Phys. 115, 10265 (2001).

bG. Auböck, J. Nagl, C. Callegari, and W. E. Ernst, J. Phys. Chem. A 111, 7404 (2007).